OpenStudy (qtpi3.14):

4th and 10th terms of a geometric sequence are 1/8 and 1/128 respectively. Find the 5th term

OpenStudy (qtpi3.14):

@ganeshie8 @Whitemonsterbunny17 @samhang @.Sam. @EclipsedStar @sweetburger

OpenStudy (mecharv):

You know the GP formula right?

OpenStudy (qtpi3.14):

whats the gp formula?

OpenStudy (mecharv):

\[a _{n}=a _{1}r ^{n-1} \] where r is common ration, a1 is the first term of the gp, n is the number of terms

OpenStudy (qtpi3.14):

What do we do now?

OpenStudy (mecharv):

We dont know a1 nor r, so we have to find them. Now you use the given value and put it in the formula, you will get two equations, using which you can eliminate a1 to find r.

OpenStudy (qtpi3.14):

\[a_4=a_1r^{4-1}=\frac{1}{8} \\ \\ a_{10}=a_1r^{10-9}=\frac{1}{128}\] is this it?

OpenStudy (mecharv):

check your a10 again?

OpenStudy (qtpi3.14):

oops

OpenStudy (mecharv):

@qtpi3.14 and can you also check the question itself? I feel there might be some mistake... just check and if it is correct we will proceed with the same.

OpenStudy (qtpi3.14):

not sure

OpenStudy (qtpi3.14):

i got \[a_1=128^2\]

OpenStudy (mecharv):

How did you get that? :O No that is wrong.

OpenStudy (qtpi3.14):

\[a_5=128^2(^6\sqrt{\frac{1}{128})}^4\]

OpenStudy (mecharv):

how did you find a1?

OpenStudy (qtpi3.14):

um I made a mistake

OpenStudy (mecharv):

Can you check your mistake?

OpenStudy (qtpi3.14):

hold on

OpenStudy (qtpi3.14):

from a4 i'm able to find a1=1/8r^3

OpenStudy (qtpi3.14):

once a1 is 1/8r^3

OpenStudy (qtpi3.14):

now substituting a1 into a10 \[a_{10}=(\frac{1}{8r^3})(r^9)=\frac{1}{128}\]

OpenStudy (mecharv):

Yes that is correct

OpenStudy (qtpi3.14):

\[r=~^6\sqrt{\frac{1}{16}}\]

OpenStudy (mecharv):

Perfect, that is correct.

OpenStudy (qtpi3.14):

\[8a_1=\frac{1}{[(\frac{1}{16})^{\frac{1}{6}}]^3}\]

OpenStudy (mecharv):

where did 8a1 come from?

OpenStudy (qtpi3.14):

from a4

OpenStudy (qtpi3.14):

i got a1=1/2

OpenStudy (mecharv):

Yeah now you know r,a1,n. Substitute in the general formula and that is the answer

OpenStudy (qtpi3.14):

thank you

OpenStudy (mecharv):

Anytime :)

OpenStudy (mecharv):

@qtpi3.14 still here? lol